OpenStudy (anonymous):

solve this de. it says to show all step in your working out...i'm not sure what i have done is 'all steps'

OpenStudy (anonymous):

\[x^3\frac{ dy }{ dx }+3x^2y=e^x\]

OpenStudy (anonymous):

so i know the answer but what would it mean by 'all steps'

OpenStudy (anonymous):

All i really did was find I(x) then solved I(x)y=integral of I(x)Q(x)

OpenStudy (anonymous):

then solved for y

OpenStudy (anonymous):

would someone share me what they would do in reference to all steps..

zepdrix (zepdrix):

What is I(x)? Is that your integrating factor that you found?

OpenStudy (anonymous):

yep!

zepdrix (zepdrix):

lol you had to do some work to find I(x).. not sure why you're confused about "show your work" lol XD

OpenStudy (anonymous):

alright this is what i did

OpenStudy (anonymous):

\[\frac{ dy }{ dx }+\frac{ 3 }{ x }y=\frac{ e^x }{ x^3 }\]\[I(x)=\exp(\int\limits_{?}^{?}\frac{ 3 }{ x })\]\[\exp(3\ln \left| x \right|)\]=\[x^3\]

OpenStudy (anonymous):

\[yI(x)=\int\limits_{?}^{?}e^x \] as we are integrating x^3.e^x/x^3

OpenStudy (anonymous):

that is simply e^x +C

OpenStudy (anonymous):

THEREFORE